# A strong–weak duality for the 1d long-range Ising model

### v1

@@ 3.2.2 Covariant derivative and Schwinger-Dyson equation
The covariant derivative allows us to understand if some operator is a primary or a descendant at the conformal fixed point. Interestingly, $D_{x}O$ can be a descendant of $O$ even if $O$ has no explicit dependence on the defect coordinate. An important example for us will be the case of $O=\hat{\sigma}_{3}$, for which we have
@@ 3.2.3 Symmetries
where $\alpha\in\mathbb{R}/2\pi\mathbb{Z}$. The model (3.9) is natural because $\mathcal{O}_{g}$ and $\mathcal{O}_{h}$ are the only relevant or marginal operators invariant under this $\mathrm{O}(2)$.
@@ 4 Renormalization group analysis
We shall distinguish between the model with $s=1$, in which case a free boundary CFT is perturbed by marginal primary operators, from the model with $s<1$, in which case a GFF with a compact and dimensionful target space is perturbed by composite operators that are not scaling operators. In the latter case we do not have at the moment a proof that the theory is renormalizable to all orders of the perturbative expansion, but we also found no evidence to the contrary in our computations. Although we hope to come back to this question in the future, it should be said that the UV completeness of the model is not needed for our purposes, as we are only interested in its IR behavior.
@@ 4.1 Beta functions and fixed point at $s=1$
where we have implicitly subtracted power-law divergences, and the dots stand for regular terms as $a\to 0$.
@@ 4.2 Spectrum at $s=1$
We can identify the first line with $\chi\sim\varphi^{3}$ (in hindsight, this is why we defined $\chi=\frac{\mathrm{i}}{\sqrt{2}}\partial\phi$). The second and third line are two linearly independent marginal operators which are uncharged under both $\mathbb{Z}_{2}$ and parity. Finally, the last line is needed to recombine with the trivial $\Delta=0$ conformal multipliet of $\sigma$ to form a conformal multiplet with $\Delta_{\sigma}>0$.
@@ 4.3 Beta functions and fixed point at $s<1$
This is yet not the final answer for the physical RG. Indeed, as explained in Section 3.2.1, partition function and correlators only depend on the combination $b_{0}-\sqrt{2}h$, hence the beta function entering the Callan-Symanzik equation is that for such a combination. In standard RG language, what this means is that while both $h$ and $b_{0}$ flow, one of them is a redundant coupling, only the combination $b_{0}-\sqrt{2}h$ being an essential coupling. Therefore, we define the physical coupling in cutoff units as $b=(b_{0}-\sqrt{2}h)a^{\delta/2}$, and obtain the beta functions
@@
namely the same as (4.13) plus the classical contribution due to the dimension of $b^{2}$. We have thus shown how a perturbative computation in the model (3.9) fully reproduces the Kosterlitz beta functions (2.21), with $\mathcal{J}=b^{2}$, plus a next-to-leading correction of order $g^{3}$.
Fixed points
@@ 4.4.1 Leading near-marginal operators at the fixed point
By standard RG arguments, linearizing the beta functions around the IR fixed point we find the linear combination of $\mathcal{O}_{h}$ and $\mathcal{O}_{g}$ that behave as scaling operators at the fixed point, along with their IR scaling dimensions $\Delta_{\pm}=1+\omega_{\pm}$. The quantities $\omega_{\pm}$ are the eigenvalues of the stability matrix:
@@ 4.4.2 Anomalous dimensions of $\hat{\sigma}_{3}$ and $\chi$
The anomalous dimension of an operator $\mathcal{O}$ can be computed by treating $\mathcal{O}$ as a perturbation with coupling $\lambda$, and computing the beta function $\beta_{\lambda}$ for this new coupling. From there, the scaling dimension is obtained from $\partial_{\lambda}\beta_{\lambda}$ evaluated at the fixed point. Considering the expectation value of $\sigma\equiv a^{-\Delta_{\sigma}}\hat{\sigma}_{3}$, where the factor of $a$ ensures that the engineering and scaling dimension match, the first non-trivial contribution is at quadratic order in $g$ and $h$, for which we obtain:
$a^{\Delta_{\sigma}}\langle\sigma(\infty)\rangle_{\mathcal{D}}$ $=\frac{\lambda}{2}\int{\rm d}y\operatorname{tr}\langle\hat{\sigma}_{3}(\infty)\hat{\sigma}_{3}(y)\rangle_{0}$
@@
from which we find the following beta function:
$\beta_{\lambda_{\sigma}}=-\frac{d\lambda}{d\log a}=2\lambda g^{2}+\dots\>.$ (4.35)
The corresponding anomalous dimension is
$\gamma_{\sigma}=\partial_{\lambda}\beta_{\lambda}=2g^{2}\quad\to\quad\gamma_{\sigma}|_{g=g_{*}}=\frac{\delta}{2}\>.$ (4.36)
That this dimension does not receive higher-order corrections in $\delta$ is guaranteed by the Schwinger-Dyson equation (3.30), which fixes $D_{x}\hat{\sigma}_{3}$ to be a descendant operator of protected scaling dimension $1+\delta/2$ at the non-trivial IR fixed point.
@@ 5.2 The crossing equations of light primaries up to $O(\sqrt{\delta})$
as well as the statement that $c_{\sigma\sigma\mathcal{P}}^{(1)}=0$ unless $\mathcal{P}=\mathcal{O}_{\pm}$. We can now also update (5.5), (5.9), and (5.13) using the correct value of $\theta$
@@ 5.3.1 Analytic functionals
$\omega^{m_{1},m_{2},m_{3},m_{4}}_{n}[f]=\int\limits_{\frac{1}{2}-i\infty}^{\frac{1}{2}-i\infty}G^{1-m_{1},1-m_{2},1-m_{3},1-m_{4}}_{1-n}(z)f(z)\frac{dz}{2\pi i}\,.$ (5.22)
@@ 5.3.4 Crossing of $\langle\sigma\chi\sigma\chi\rangle$ at $O(\delta)$
To go to the second line, we used a standard transformation of $_{2}F_{1}$ hypergeometric functions. To go to the third line, we used (5.52) and the positivity of coefficients of the $y$ expansion in the $_{2}F_{1}$. To go to the last line, we explicitly evaluated the sum on the third line and took the limit $|y|\rightarrow 1$.
@@ 5.3.6 Crossing of $\langle\sigma\sigma\sigma\chi\rangle$ at $O(\delta^{3/2})$
for all $n\geq 0$. Note that based on previous information, we could have predicted that $A^{\sigma\sigma\sigma\chi}_{n}=0$ for $n$ even. This is because $c^{(2)}_{\sigma\sigma m}\neq 0$ only if $(-1)^{J_{m}}=1$, and $c^{(1)}_{\sigma\chi m}\neq 0$ only if $(-1)^{J_{m}}=(-1)^{\Delta^{(0)}_{m}-1}$. For $n$ odd, (5.70) is a new fact.
@@ 6 Conclusions and future directions
Regarding the fact that $\mathcal{O}_{g}$ is not a scaling operator of the unperturbed theory, we make the following remarks. Firstly, even in standard massive perturbation theory, operators such as $\varphi^{4}$ are not scaling operators, they only become such in the UV limit, when the mass is neglected. Similarly, one could consider the UV limit of our model, in which case $b_{0}\to 0$. In this limit, $\mathcal{O}_{g}$ reduces to a scaling operator of vanishing dimension. Secondly, in that same limit, one could expand $\mathcal{O}_{g}$ in an infinite series of relevant scaling operators. Since the $U(1)$ symmetry protects their relative coefficients, they still only give rise to a single coupling. Unfortunately, we have not obtained a proof of renormalizability to all orders in this framework. Having said that, we found no evidence that counterterms of irrelevant operators should be added, up to second order of the perturbative expansion. This also leads to our third remark: even if the model (3.9) turned out to not be UV complete (which we doubt), this would not affect its usefulness for the construction of the IR CFT. In fact, we are not proposing this model as a new fundamental theory, but rather as a tool to construct an IR CFT via standard RG computations. In the spirit of Wilsonian RG, the presence of irrelevant operators in the bare action has no effect on low-energy physics: the modern view is actually that all QFTs are only effective field theories. Lastly, the RG computations of Section 4, and their agreement with the bootstrap computations of Section 5, corroborate our point of view, showing that the model provides a perturbative computational framework for extracting CFT data of the 1d LRI near $s=1$.
@@ Appendix A Generalized free field on the line
A $d$-dimensional generalized free field theory (GFF) is the quantum field theory of a single scalar $\varphi$, whose only nonvanishing connected $n$-point function is the two-point function $\langle\varphi(x)\varphi(y)\rangle\equiv C(x-y)$, assumed to be invariant under rotations and translations. All other (not connected) $n$-point functions are obtained as sums over Wick contractions. In other words, a GFF is a centered Gaussian measure, and as such it can be defined without introducing any action functional, see e.g. [61, 101].
@@
We restrict now to a conformal GFF in $d=1$. For $\Delta_{\varphi}>0$, it provides the simplest example of a 1d CFT as defined in Section 2.2.1. For $\Delta_{\varphi}<0$, its definition and conformal properties are slightly more subtle, but it is also a well studied random field, associated to the fractional Brownian motion with so-called Hurst parameter $H=-\Delta_{\varphi}$, see for example [57, 58, 103]. The intermediate case $\Delta_{\varphi}=0$ corresponds to the log-correlated Gaussian field [104], which shares many technical aspects with the $\Delta_{\varphi}<0$ case, and which in $d=1$ coincides with the usual 2d free field theory restricted to a line, as we will recall below.

### v2

@@ 3.2.2 Covariant derivative and Schwinger-Dyson equation
The covariant derivative allows us to understand if some operator is a primary or a descendant at the conformal fixed point. Interestingly, $D_{x}O$ can be non-zero even if $O$ has no explicit dependence on the defect coordinate. An important example for us will be the case of $O=\hat{\sigma}_{3}$, for which we have
@@ 3.2.3 Symmetries
where $\alpha\in\mathbb{R}/2\pi\mathbb{Z}$. The model (3.9) is natural, in the sense that at $s=1$ the operators $\mathcal{O}_{g}$ and $\mathcal{O}_{h}$ are the only relevant or marginal operators invariant under this $\mathrm{O}(2)$.
@@ 4 Renormalization group analysis
We shall distinguish between the model with $s=1$, in which case a free boundary CFT is perturbed by marginal primary operators, from the model with $s<1$, in which case a GFF with a compact and dimensionful target space is perturbed by composite operators that are not scaling operators. In the latter case we do not have at the moment a proof that the theory is renormalizable to all orders of the perturbative expansion, but we also found no evidence to the contrary in our computations. One reason for expecting (3.9) to be renormalizable is that, as we will see below, at $\delta>0$ the theory has a UV fixed point, where both $\mathcal{O}_{g}$ and $\mathcal{O}_{h}$ become relevant scaling operators. Therefore, at least in principle, we could start from that manifestly renormalizable setting in order to construct our flow. In practice, this is not very useful, because in this case $\mathcal{O}_{g}$ is strictly relevant (i.e. super-renormalizable) and the UV fixed point turns out to be far from the IR one. [^20] There is a similar problem when deriving the BKT flow diagram from the sine-Gordon model [62]. Moreover, at $\delta=0$ one of the relevant perturbations of such UV fixed point becomes exactly marginal, and it becomes impossible to reach the IR fixed point of interest in this way. The solution is to work with (3.9) away from the UV fixed point at $\delta>0$, but perturbatively in $\delta$ and thus close to the fixed point of the $\delta=0$ model. Therefore, we begin our analysis from the $\delta=0$ case.
@@ 4.1 Beta functions and fixed point at $s=1$
where we have implicitly subtracted power-law divergences, and the dots stand for vanishing terms as $a\to 0$.
@@ 4.2 Spectrum at $s=1$
We can identify the first line with $\chi\sim\varphi^{3}$ (in hindsight, this is why we defined $\chi=\frac{\mathrm{i}}{\sqrt{2}}\partial\phi$ in (3.10)). The second and third line are two linearly independent marginal operators which are uncharged under both $\mathbb{Z}_{2}$ and parity. Finally, the last line is needed to recombine with the trivial $\Delta=0$ conformal multipliet of $\sigma$ to form a conformal multiplet with $\Delta_{\sigma}>0$.
@@ 4.3 Beta functions and fixed point at $s<1$
This is yet not the final answer for the physical RG. Indeed, as explained in Section 3.2.1, partition function and correlators of $U(1)$-singlet operators only depend on the combination $b_{0}-\sqrt{2}h$, hence the beta function entering the Callan-Symanzik equation for correlators of singlet oprators is that for such a combination. In standard RG language, what this means is that while both $h$ and $b_{0}$ flow, one of them is a redundant coupling, only the combination $b_{0}-\sqrt{2}h$ being an essential coupling. Therefore, we define the physical coupling in cutoff units as $b=(b_{0}-\sqrt{2}h)a^{\delta/2}$, and obtain the beta functions
@@
namely the same as (4.13) plus the classical contribution due to the dimension of $b^{2}$. We have thus shown how a perturbative computation in the model (3.9) fully reproduces the Kosterlitz beta functions (2.21), and thus the flow of figure 3, with $\mathcal{J}=b^{2}$, plus a next-to-leading correction of order $g^{3}$.
4.3.1 Fixed points
@@
IR fixed points and gauging.
The fixed point (4.22) is the one that corresponds to the 1d CFT for the critical LRI. We should stress that the relevant fixed points are those of $\beta_{b^{2}}$, and not those of $\beta_{h}$, for the reasons explained above. In other words, we are seeking fixed points for the singlet sector of the model (3.9), that is, the gauged model. This remark is non-trivial because gauging and seeking for fixed points do not commute: the condition $\beta_{b_{0}}=\beta_{h}=0$ is stronger than $\beta_{b^{2}}=0$, and its only fixed point that is perturbatively accessible is the trivial one at $b_{0}=h=g=0$; conversely, at the nontrivial fixed point of $\beta_{b^{2}}$, the dimensionless couplings $b_{0}a^{\delta/2}$ and $ha^{\delta/2}$ keep flowing, hence we must restrict to the $U(1)$-singlet sector in order to have a CFT, because correlators of non-singlet operators in general depend separately on $b_{0}$ and $h$.
UV fixed point.
Besides the nontrivial fixed point (4.22), the beta functions (4.21) admit also a trivial fixed point at $g=b^{2}=0$. At $\delta>0$, both $g$ and $b^{2}$ are relevant perturbations of this fixed point, which thus is a UV fixed point, as it is also clear from figure 3. Such UV fixed point describes a a noncompact GFF $\phi$ plus a decoupled qubit. Around this point, the model (3.9) corresponds to a standard perturbation by relevant operators. Notice that the operator $\mathcal{O}_{g}$ reduces to $\hat{\sigma}_{+}+\hat{\sigma}_{-}$, which is still nontrivial because it does not commute with $\mathcal{O}_{h}$: we can repeat the analysis of section 3.2.1, and find that the effective $b^{2}$ increases under the effect of the $\mathcal{O}_{h}$ term. Therefore, we see that from an RG perspective we can qualitatively view $b^{2}$ as the analogue of a mass parameter $m^{2}$: perturbing a standard massless free boson $\varphi$ by a $\varphi^{2}$ term we generate a mass, and an RG flow; if instead we start directly with a massive boson, self-interactions such as $\varphi^{4}$ are not scaling operators, they are such only in the UV limit.
Absence of phase transition at $s>1$.
For $\delta<0$, i.e. $s>1$, the non-trivial fixed point (4.22) becomes complex (real $b^{2}_{*}$, imaginary $g_{*}$). In this case the flow in the real $\{g,b^{2}\}$ plane displays only the trivial fixed point at $g=b^{2}=0$, which now has one irrelevant ($b^{2}$) and one relevant ($g$) perturbation. Therefore, trajectories always flow towards large $g$, i.e. the symmetric phase, unless we start exactly from $g=0$ (i.e. zero temperature), in which case we end up at the trivial fixed point, where the qubit degree of freedom is decoupled from the GFF. In other words, the physics of the 1d SRI is reproduced at $s>1$.
@@ 4.4.1 Leading near-marginal operators at the fixed point
By standard RG arguments, linearizing the beta functions around the IR fixed point (4.22) we find the linear combination of $\mathcal{O}_{h}$ and $\mathcal{O}_{g}$ that behave as scaling operators at the fixed point, along with their IR scaling dimensions $\Delta_{\pm}=1+\omega_{\pm}$. The quantities $\omega_{\pm}$ are the eigenvalues of the stability matrix:
@@ 4.4.2 Anomalous dimensions of $\hat{\sigma}_{3}$ and $\chi$
The anomalous dimension of an operator $\mathcal{O}$ can be computed by treating $\mathcal{O}$ as a perturbation with coupling $\lambda$, and computing the beta function $\beta_{\lambda}$ for this new coupling. From there, the scaling dimension is obtained from $\partial_{\lambda}\beta_{\lambda}$ evaluated at the fixed point at $\lambda=0$, via the formula $\partial_{\lambda}\beta_{\lambda}|_{g*;\lambda=0}=\Delta_{\mathcal{O}}-d$. Considering the expectation value of $\hat{\sigma}_{3}$, the first non-trivial contribution is at quadratic order in $g$ and $h$, for which we obtain:
$\langle\hat{\sigma}_{3}(\infty)\rangle_{\mathcal{D}}$ $=\frac{\lambda}{2}\int{\rm d}y\operatorname{tr}\langle\hat{\sigma}_{3}(\infty)\hat{\sigma}_{3}(y)\rangle_{0}$
@@
and introducing the dimensionless coupling $\tilde{\lambda}=a\lambda$, we find the following beta function:
$\beta_{\lambda_{\sigma}}=-\frac{d\tilde{\lambda}}{d\log a}=-\tilde{\lambda}+2\tilde{\lambda}g^{2}+\dots\>.$ (4.35)
Given that the canonical dimension of $\hat{\sigma}_{3}$ is zero, we can identify $\Delta_{\sigma}$ with the anomalous dimension, which is thus
$\gamma_{\sigma}=1+\partial_{\lambda}\beta_{\lambda}=2g^{2}\quad\to\quad\Delta_{\sigma}=\gamma_{\sigma}|_{g=g_{*}}=\frac{\delta}{2}\>.$ (4.36)
That this dimension does not receive higher-order corrections in $\delta$ is guaranteed by the Schwinger-Dyson equation (3.30), which fixes $D_{x}\hat{\sigma}_{3}$ to be a descendant operator of protected scaling dimension $1+\delta/2$ at the non-trivial IR fixed point. In order to match the engineering and scaling dimension of the spin operator, we define $\sigma\equiv a^{-\Delta_{\sigma}}\hat{\sigma}_{3}$, that we will use in the following.
@@ 5.2 The crossing equations of light primaries up to $O(\sqrt{\delta})$
as well as the statement that $c_{\sigma\sigma\mathcal{P}}^{(1)}c_{ab\mathcal{P}}^{(0)}=0$ unless $\mathcal{P}=\mathcal{O}_{\pm}$. We can now also update (5.5), (5.9), and (5.13) using the correct value of $\theta$
@@ 5.3.1 Analytic functionals
$\omega^{m_{1},m_{2},m_{3},m_{4}}_{n}[f]=\int\limits_{\frac{1}{2}-i\infty}^{\frac{1}{2}+i\infty}G^{1-m_{1},1-m_{2},1-m_{3},1-m_{4}}_{1-n}(z)f(z)\frac{dz}{2\pi i}\,.$ (5.22)
@@ 5.3.2 Crossing of $\langle\sigma\sigma\sigma\sigma\rangle$ at $O(\delta^{2})$
Incidentally, we should ask whether there could have been $\Delta\geq 2$ contributions to $\langle\sigma\sigma\sigma\sigma\rangle$ already at $O(\delta)$. In fact, it is not hard to see that such contributions are forbidden. The crucial point is that, as we remarked in Section 5.2.1, the total contribution of $\mathds{1}$, $\mathcal{O}_{+}$, and $\mathcal{O}_{-}$ to $\langle\sigma\sigma\sigma\sigma\rangle$ at $O(\delta)$ is already crossing symmetric on its own. We can then use the analytic functionals $\omega^{0,0,0,0}_{n}$ to solve for the sum of $(c^{(1)}_{\sigma\sigma\mathcal{P}})^{2}$ over all $\mathcal{P}$ with $\Delta^{(0)}_{\mathcal{P}}=n\geq 2$, analogously to (5.36). The difference is that at $O(\delta)$, $\mathcal{G}^{s,L}_{\sigma\sigma\sigma\sigma}(1-z)-\mathcal{G}^{s,L}_{\sigma\sigma\sigma\sigma}(z)=0$, and applying the analytic functionals gives vanishing total OPE squared.
@@ 5.3.4 Crossing of $\langle\sigma\chi\sigma\chi\rangle$ at $O(\delta)$
To go to the second line, we used a standard transformation of $_{2}F_{1}$ hypergeometric functions. To go to the third line, we used (5.52) and the positivity of coefficients of the series expansion in the $_{2}F_{1}$. To go to the last line, we explicitly evaluated the sum on the third line and took the limit $z\rightarrow\infty$.
@@ 5.3.6 Crossing of $\langle\sigma\sigma\sigma\chi\rangle$ at $O(\delta^{3/2})$
for all $n\geq 0$. The equation holds because all of the hitherto unknown parameters only appear through an overall prefactor, multiplying the function (2z1)/z(2z-1)/z. The latter function satisfies (5.70). Note that based on previous information, we could have predicted that $A^{\sigma\sigma\sigma\chi}_{n}=0$ for $n$ even. This is because $c^{(2)}_{\sigma\sigma m}\neq 0$ only if $(-1)^{J_{m}}=1$, and $c^{(1)}_{\sigma\chi m}\neq 0$ only if $(-1)^{J_{m}}=(-1)^{\Delta^{(0)}_{m}-1}$. For $n$ odd, (5.70) is a new fact.
@@ 6 Conclusions and future directions
Regarding the fact that $\mathcal{O}_{g}$ is not a scaling operator of the unperturbed theory, we make the following remarks. Firstly, even in standard massive perturbation theory, operators such as $\varphi^{4}$ are not scaling operators, they only become such in the UV limit, when the mass is neglected. Similarly, one could consider the UV limit of our model, in which case $b_{0}\to 0$. In this limit, $\mathcal{O}_{g}$ reduces to a scaling operator of vanishing dimension. Moreover, in that same limit, one could expand $\mathcal{O}_{g}$ in an infinite series of relevant scaling operators. Since the $U(1)$ symmetry protects their relative coefficients, they still only give rise to a single coupling. Unfortunately, we have not obtained a proof of renormalizability to all orders in this framework. Having said that, we found no evidence that counterterms of irrelevant operators should be added, up to second order of the perturbative expansion. Lastly, the RG computations of Section 4, and their agreement with the bootstrap computations of Section 5, corroborate our point of view, showing that the model provides a perturbative computational framework for extracting CFT data of the 1d LRI near $s=1$.
@@
It is interesting to compare the perturbative RG and the conformal bootstrap treatment. The RG treatment provides a self-contained definition of the model, and a framework where any quantity can in principle be calculated to any order in perturbation theory. In practice, the conformal bootstrap arrives at the final results more efficiently, by virtue of operating directly at the conformal fixed point, and by focusing only on the pertinent quantities. Note that the only input from the RG model needed to perform the conformal bootstrap calculation was the knowledge of the spectrum of scaling dimensions strictly at $\delta=0$.
One may wonder whether the conformal bootstrap could be used to completely replace the perturbative RG treatment. Perturbative RG can, at least in principle, determine all the CFT data to all orders in $\sqrt{\delta}$. We have seen that, to the first few orders, the conformal bootstrap does determine at least some CFT data with greater ease. To continue in this manner to higher orders and for other pieces of the CFT data, it will be necessary to enlarge the set of external operators in the four-point functions whose crossing symmetry is imposed. It is plausible that crossing symmetry of all four-point functions does determine the CFT data to all orders (and even non-perturbatively at finite $\delta$). However, we are not aware of any result that would imply that in order to determine a given finite set of CFT data to a given finite perturbative order, only a finite number of crossing equations will be needed. In this sense, perturbative RG has not been superseded. However, see [82] for an analogous situation in geometry where the bootstrap provably determines all the data.
@@ Appendix A Generalized free field on the line
A $d$-dimensional generalized free field theory (GFF) is the quantum field theory of a single scalar $\varphi$, whose only nonvanishing connected $n$-point function is the two-point function $\langle\varphi(x)\varphi(y)\rangle\equiv C(x-y)$, assumed to be invariant under rotations and translations. All other (not connected) $n$-point functions are obtained as sums over Wick contractions. In other words, a GFF is a centered Gaussian measure, and as such it can be defined without introducing any action functional, see e.g. [64, 105], whose explicit expression we postpone to the following subsections.
@@
We restrict now to a GFF in $d=1$, reserving the name GFF precisely for the case of a fractional Gaussian field. For $\Delta_{\varphi}>0$, it provides the simplest example of a 1d CFT as defined in Section 2.2.1. For $\Delta_{\varphi}<0$, its definition and conformal properties are slightly more subtle, but it is also a well studied random field, associated to the fractional Brownian motion with so-called Hurst parameter $H=-\Delta_{\varphi}$, see for example [58, 59, 107]. The intermediate case $\Delta_{\varphi}=0$ corresponds to the log-correlated Gaussian field [108], which shares many technical aspects with the $\Delta_{\varphi}<0$ case, and which in $d=1$ coincides with the usual 2d free field theory restricted to a line, as we will recall below.